# Field-based Molecule Generation

### v1

@@ Appendix B Field parameters
Appendix C Model and optimization parameters
@@
We use 100 timesteps $t\in[0,1]$ for our diffusion processes, the $\boldsymbol{\epsilon}$ parametrization, and the modified cosine scheduler for $\alpha_{t}$, as presented in main Section 3. The conditional scale for the classifier-free guidance was kept to $\beta=2$ during the generation phase, for our main results (unless explicitly mentioned otherwise).
@@
We optimize the $\mathcal{L}_{\mathrm{simple}}$ objective using Adam optimizer with $8\cdot 10^{-5}$ learning rate and (0.9, 0.99) for the Adam’s $\beta$ parameters. The models are trained using batch sizes of 32 and 64, for GEOM and QM9 datasets, respectively.
@@
We use a three-dimensional adaptation of the U-Net employed by (Ho et al., 2020b). For both GEOM and QM9 datasets, we use the same hyperparameters, except for the batch size. The architecture consists of three U-Net layers, with linear attention for each resolution level, except the last one, which uses full attention. Its initial convolution converts the 8 and 11 channels (for QM9 and GEOM datasets, respectively), into a hidden representation of $\mathbb{R}^{128\times H\times W\times D}$, which is halved for each of the $H,W,D$ dimensions, while its channel size is multiplied by 1, 2, and 3.
@@
For the QM9 experiments, we use four A100 GPUs (40GB memory version) for 180 hours, and 1.2 million iterations (or about 780 epochs). On the GEOM dataset, we used the same four A100 GPUs for 330 hours and trained our explicit H GEOM model for 1.32 million iterations (6 epochs), and our implicit one for 1.2 million iterations (5.5 epochs).
Appendix D Molecule graph extraction parameters and algorithms
@@
The algorithm works by defining a simple, recursive flood-fill algorithm for retrieving the neighbors. Assuming the RBF functions have a small $\sigma$ parameter and reasonably small amounts of noise produced by the diffusion model, it will never fail to extract the correct number of atoms and their approximate positions from the fields.
@@
The initial atom positions retrieved in the $A_{\mathbf{m}}$ sets, are further optimized using Equation 13. Note how both atom numbers and their types are obtained using the peak extraction algorithm, and the optimization only fine-tunes their positions.
Extracting bonds
The distances between any two bonds $b_{ij}$ and $b_{jk}$ are half as large as the distances between the atoms $a_{i}$ and $a_{k}$ ($||\mathbf{m}_{ij}-\mathbf{m}_{jk}||_{2}=||\mathbf{m}_{i}-\mathbf{m}_{k}||_{2}/2$). This may create a significantly larger overlap between the bond RBF components, compared to atoms. Therefore, we treat the bond assignment differently and do not reuse the same peak extraction algorithm. Using the optimized atom positions $\mathbf{M}_{a}$, we check for all the atom pairs $\mathbf{m}_{i},\mathbf{m}_{j}\in\mathbf{M}_{a}$ that are within some predetermined threshold distance $||\mathbf{m}_{i}-\mathbf{m}_{j}||_{2}<d_{1}$ and determine $B_{ij,b}=\{u_{b}(\mathbf{x});||\mathbf{x}-\mathbf{m}_{ij}||_{2}<d_{2},u_{b}(\mathbf{x})>r_{2}\}$. A covalent bond type $b$ is added between atoms $i$ and $j$ if $B_{ij,b}\neq\varnothing$.
To ensure we don’t miss bonds due to the low resolution, we select a large sphere (determined by its radius $d_{2}$) around the approximate bond positions $\mathbf{m}_{ij}$, within which we check for values greater than $r$. Additionally, since the true $\mathbf{m}_{ij}$ positions may land at the edge of a voxel, the maximum value corresponding to bonds $b_{ij}$ in the tensors $\mathbf{u}_{b}$ can be small, and therefore we chose a low threshold $r$. This may then recover bonds at positions where the model did not generate a true RBF bond component.
Optimizing $\gamma_{ij}$
To remove erroneous bonds, extracted due to our lenient bond-extraction parameters, we optimize the $\gamma_{ij}$ assigned to each bond $b_{ij}$, assuming the mean bond positions $\mathbf{m}_{ij}$ fixed. Let $\gamma_{b}=\{\gamma_{ij};b_{ij}=b\}$ (where $b_{ij}$ is initially assigned based on the condition $B_{ij,b}\neq\varnothing$), then optimize the RBF components $\gamma_{b}$ using equation 14. In Figure 9, we analyze how robust the method is to various degrees of uniform noise. We choose 300 molecules from the QM9 dataset and gradually add more and more noise to the voxelized fields $\mathbf{u}$.
@@
We choose threshold parameter $t=0.3$ in Algorithm 1. For bond detection, we choose a large distance $d_{1}$ of $0.35$Å (for reference, the margin used to detect bond type 1 based on distance in the ablation Table 5 is $0.1$Å). The distance $d_{2}=0.45$ will result in the method checking for the closest midpoint voxel to $\mathbf{m}_{ij}$ and a few others directly in contact with it, while the bond threshold is set to $r=0.3$. Both the mean positions $\mathbf{m}$ and bond RBF component factors $\gamma_{ij}$ are optimized using gradient descent for 500 iterations and a learning rate of $5$. All parameters were selected according to robustness analyses, such as the one illustrated in Figure 9, based on the training data.
Appendix E Generated samples
Figure 10: Random valid samples from FMG trained on QM9.
Figure 11: Random valid samples from FMG trained on GEOM.
Appendix F Classificaiton guidance analysis
We determine how accurately the classification free guidance results in molecules of the actual desired number of atoms. We compute the accuracy as the percentage of generated molecules in the correct conditioning bin $N$, the L1 distance between the binned number of atoms generated by the model $\tilde{N}$, and the conditioning bin $N$ as $\min(|\tilde{N}_{\text{high}}-N_{\text{low}}|,|N_{\text{high}}-\tilde{N}_{\text{low}}|)$ (or 0, when $N=\tilde{N}$), and “L1 lower” is the fraction of the total L1 distance attributed to incorrectly generated bins because of a strictly lower number of atoms generated $\tilde{N}<N$ (as a measure of a left or right shift of the distribution).
@@ Appendix F Classificaiton guidance analysis
Similar results in terms of accuracy and L1 distance can be noted in the GEOM dataset (relative to the much larger possible number of atoms), but now the shift is in the opposite direction. This could be an artifact of the number of training iterations: early on in the QM9 dataset training experiments, we observed that the model tends to generate a lower number of atoms and bonds, which likely explains this opposite shift of the GEOM method. For reference, our GEOM model has been trained for about 6 epochs, compared to its QM9 counterpart (about 780).
Figure 12: Histogram of the generated number of atom classes $\tilde{N}$ and the conditioning used $N$. Small and more significant atom number distribution shifts occur for QM9 and GEOM datasets, respectively. Together with the result of Table 5 which shows how critical it is to enforce the correct number of atoms for TVa, this result suggests that our method may remove atoms for stability, inducing an atom distribution shift.
Appendix G Visualization
Figure 3 is a simplified illustration meant to conceptually show the noise scheduler’s effect, and it is perhaps not accurately representing an actual visualization of FMG generating a molecule (please refer to Figure 13 for such an illustration).
Figure 13: Three-dimensional visualization of a generated molecule by FMG on QM9. The upper and lower images depict the summed values (normalized to $[0,1]$ at each step) over all bond $\mathbf{u}_{b,1}+\mathbf{u}_{b,2}+\mathbf{u}_{b,3}$ and atom $\mathbf{u}_{\texttt{H}}+\mathbf{u}_{\texttt{C}}+\mathbf{u}_{\texttt{O}}+\mathbf{u}_{\texttt{N}}+\mathbf{u}_{\texttt{F}}$ fields, respectively. The same effect as in the main paper Figure 3 can be seen, where bonds become visible through the noise quicker than the atoms. Circles and lines are added for reference, denoting atoms and bonds extracted at the last time step.
Figure 3 of the main text is done using the maximum $L_{2}$ loss between noisy $\mathbf{u}_{t}$ (for various $t\in[0,1]$) and the determined RBF function at step $t=0$. Using a similar loss function as in equation 13, we determine the fidelity of each timestep atom and bond fields ($\mathbf{u}_{b,t}$, $\mathbf{u}_{b,a}$) for atom $a$ and bond $b$ channels separately, to the final extracted RBF components based on $\mathbf{u}_{0}$. Using the maximum loss values over each channel separately, we vary the illustration’s transparency parameter for various bond and atom types over time $t$.
@@ Appendix H Performance metric details
Percentage of valid molecules, computed using RDKit’s SanitizeMol function.
@@
Novelty:
percentage of generated molecular graphs that are not present in the training data (after canonicalization).
@@
$\text{TV}_{\text{a}}=\sum_{a\in\mathcal{A}}\sum_{N_{a}=i}^{N}|p_{d}(N_{a})-p_{g}(N_{a})|$
$\text{TV}_{\text{b}}=\sum_{b\in\mathcal{B}}\sum_{N_{b}=i}^{N}|p_{d}(N_{b})-p_{g}(N_{b})|,$
where $N_{a},N_{b}$ are the numbers of atoms and bonds of type $b$, $a$ and $b$ are the atom and bond types, $\mathcal{A}$ and $\mathcal{B}$ are the sets of all unique atoms and bonds, $p_{d}$ and $p_{g}$ are probability distributions determined on data and generated statistics, and ||$\cdot$|| ⋅ | is the absolute value.
@@
BL and BA:
We determine the Wasserstein W1 distance between the generated and data bond angles and bond lengths.
The angle distributions are defined, for each atom type, by all atoms that have two or more bonds, and we consider all pair-wise angles between position vectors $\mathbf{m}_{i}-\mathbf{m}_{k}$, and $\mathbf{m}_{j}-\mathbf{m}_{k}$, for any two atoms $i$ and $j$ connected to the same central atom $k$. We consider the average W1 distance over all atom types, and we denote it by BA.
@@
We use discretized bins of length $l_{\theta}=0.1$ degrees for angles, and $l_{d}=0.01$Å for bond Euclidean distances, and approximate the generated and data cumulative distribution functions $F_{g}$ and $F_{d}$ and compute:
@@
$\text{BL}=l_{b}\frac{1}{N_{\mathcal{B}}}\sum_{b\in\mathcal{B}}\sum_{x\in B_{b}}|F_{d,b}(x)-F_{g,b}(x)|,$
where the bin upper limits $B_{a}$ and $B_{b}$ are chosen to contain all maximum and minimum values of the generated and training data samples, $N_{\mathcal{A}}$ and $N_{\mathcal{B}}$ are the number of unique atoms and bonds considered (e.g., we excluded H and F for BA computations, as these atoms never have two bonds).

### v2

@@ Appendix C Field parameters
The field hyperparameters were selected during early experimentation. We focused on the effect of field hyperparameters on fitting $u_{a}(x)$ and $u_{b}(x)$ to the generated fields $\boldsymbol{u}_{a}$ and $\boldsymbol{u}_{b}$ (Equations 13 and  14 and Appendix E). We compute the ”extraction accuracy”, which measures how many molecules have the same molecular graph extracted from the fields they create. We additionally report the average ”distance from original”, computed as the average Euclidean distance between the correct atom positions and the ones extracted. For both, we determine how robust they are for various levels of noise applied to the fields. Ablation of field hyperparameters for the resulting diffusion model’s accuracy was not performed, due to the heavy computational resources required.
Field standard deviation $\sigma$: Intuitively, larger standard deviations $\sigma$ would facilitate smaller receptive field CNN kernels to capture longer distance neighboring RBF components. At the same time, we require that molecular graphs and conformations are correctly extracted. Figure 10 shows how increasing $\sigma$ worsens the extraction accuracy, and our choice ($\sigma^{2}=0.05$) prioritized molecular conformation (distance from original) and graph (Extraction accuracy) reliability over noise levels.
Figure 10: Noise robustness for the graph extraction procedure for various $\sigma$ field values.
Field component weight $\gamma$: The component weights may be thought of as an input normalization parameter. Therefore, we did not consider other values than $0.176$, which should normalize inputs to roughly values between $[0,1]$.
Resolution $r$: In general, it is desired to have as low resolution as possible, as this will influence the input dimensionality and, correspondingly, the depth of the parametrizing NN for limited computational resources. At the same time, the accurate extraction of conformers and graphs must be ensured. Figure 11 shows how lowering the resolution from 0.33 (our choice) to 0.44Å worsens ”extraction accuracy” and ”distance from original” across noise levels. We note, however, that $r=0.44$ may still be a valid choice, assuming the resulting NN would have low enough amounts of noise.
Figure 11: Noise robustness for the graph extraction procedure for two resolution levels.
Appendix D Model and optimization parameters
@@
We use 100 timesteps $t\in[0,1]$ for our diffusion processes, the $\bm{\epsilon}$ parametrization, and the modified cosine scheduler for $\alpha_{t}$, as presented in main Section 4. The conditional scale for the classifier-free guidance was kept to $\beta=2$ during the generation phase, for our main results (unless explicitly mentioned otherwise).
@@
We optimize the $\mathcal{L}_{\mathrm{simple}}$ objective using Adam optimizer with $8\cdot 10^{-5}$ learning rate and (0.9, 0.99) for the Adam’s $\beta$ parameters. The models are trained using batch sizes of 32 and 64, for GEOM-Drugs and QM9 datasets, respectively.
@@
We use a three-dimensional adaptation of the U-Net employed by (Ho et al., 2020). For both GEOM-Drugs and QM9 datasets, we use the same hyperparameters, except for the batch size. The architecture consists of three U-Net layers, with linear attention for each resolution level, except the last one, which uses full attention. Its initial convolution converts the 8 and 11 channels (for QM9 and GEOM-Drugs datasets, respectively), into a hidden representation of $\mathbb{R}^{128\times H\times W\times D}$, which is halved for each of the $H,W,D$ dimensions, while its channel size is multiplied by 1, 2, and 3.
@@
For the QM9 experiments, we use four A100 GPUs (40GB memory version) for 180 hours, and 1.2 million iterations (or about 780 epochs). On the GEOM-Drugs dataset, we used the same four A100 GPUs for 330 hours and trained our explicit H GEOM-Drugs model for 1.32 million iterations (6 epochs), and our implicit one for 1.2 million iterations (5.5 epochs).
Appendix E Molecule graph extraction parameters and algorithms
@@
The algorithm works by defining a simple, recursive flood-fill algorithm for retrieving the neighbors. Assuming the RBF functions have a small $\sigma$ parameter and reasonably small amounts of noise produced by the diffusion model, it should never fail to extract the correct number of atoms and their approximate positions from the fields.
@@
The initial atom positions retrieved in the $A_{\mathbf{m}}$ sets are further optimized using Equation 13. Note how both atom numbers and their types are obtained using the peak extraction algorithm, and the optimization only fine-tunes their positions $\{\mathbf{m}_{i}\}$.
Bond detection
The distances between any two bonds $b_{ij}$ and $b_{jk}$ are half as large as the distances between the atoms $a_{i}$ and $a_{k}$, meaning $||\mathbf{m}_{ij}-\mathbf{m}_{jk}||_{2}=||\mathbf{m}_{i}-\mathbf{m}_{k}||_{2}/2$. This may create a significantly larger overlap between the bond RBF components than the atoms’ RBF components. Therefore, we treat the bond assignment differently and do not reuse Algorithm 1. Using the optimized atom positions $\mathbf{M}_{a}$, we check for all the atom pairs $\mathbf{m}_{i},\mathbf{m}_{j}\in\mathbf{M}_{a}$ that are within some predetermined threshold distance $||\mathbf{m}_{i}-\mathbf{m}_{j}||_{2}<d_{1}$ and create sets
$B_{ij,b}=\{\mathbf{u}_{b}(\mathbf{x});||\mathbf{x}-\mathbf{m}_{ij}||_{2}<d_{2},\mathbf{u}_{b}(\mathbf{x})>t_{b},\mathbf{x}\in\mathcal{G}\},$
where $\mathcal{G}$ consists of the evaluated field locations, $t_{b}$ is an additional bond threshold, and $d_{2}$ is the radius of the sphere centered on $\mathbf{m}_{ij}$, around which we check for $\mathbf{u}_{b}$ values. A covalent bond type $b$ is added between atoms $i$ and $j$ if $B_{ij,b}\neq\varnothing$.
To ensure we don’t miss bonds due to the low resolution, we select a large sphere (determined by its radius $d_{2}$) around the approximate bond positions $\mathbf{m}_{ij}$ and small threshold values $t_{b}$. This may then recover bonds at positions where the model did not generate a true RBF bond component.
Bond optimization $\gamma_{ij}$
To remove erroneous bonds, extracted due to our lenient bond-extraction parameters ($d_{1},d_{2},t_{b}$), we optimize the $\gamma_{ij}$ values for all $B_{ij,b}\neq\varnothing$ assuming the mean bond positions $\mathbf{m}_{ij}$ fixed, using equation 14. In Figure 12, we analyze how robust the method is to various degrees of uniform noise. We choose 300 molecules from the QM9 dataset and gradually increase the noise amount to the voxelized field values $\mathbf{u}$.
@@
Figure 13 illustrates a bond that was initially extracted based on $B_{ij,b}\neq\varnothing$, but removed after optimizing and finding the low $\gamma_{ij}$ value.
Figure 13: Upper left subfigure illustrates a noisy covalent bond type 1 channel values $\tilde{\mathbf{u}}_{b_{1}}=\mathbf{u}_{b_{1}}+\bm{\epsilon}$ (according to the color bar), where $\bm{\epsilon}$ has the same dimension as $\mathbf{u}_{b_{1}}$ and $\bm{\epsilon}_{ijk}\sim\mathcal{U}(0,0.1)$. The bond marked with “x” on the left is removed since its optimized $\gamma_{ij}$ value (right bar plot) is significantly lower than the rest.
@@
We choose threshold parameter $t=0.3$ in Algorithm 1. For bond detection, we choose a large distance $d_{1}$ of $0.35$Å (for reference, the margin used to detect bond type 1 based on distance in the ablation Table 3 is $0.1$Å). The distance $d_{2}=0.45$ will result in the method checking for the closest midpoint voxel to $\mathbf{m}_{ij}$ and a few others directly in contact with it. The bond threshold is set to $t_{b}=0.3$. Both the mean positions $\mathbf{m}$ and bond RBF component factors $\gamma_{ij}$ are optimized using gradient descent for 500 iterations and a learning rate of $5$. All parameters were selected according to robustness analyses, such as the one illustrated in Figure 12, based on the training data.
Appendix F Generated samples
F.1 QM9 samples
Figures 14 and 15 depict a set of random valid samples from a model trained on QM9.
Figure 14: Random valid samples from FMG trained on QM9. To the left of each molecule, the top 200 largest $\mathbf{u}_{a}$ and $\mathbf{u}_{b}$ values are shown along their corresponding extracted atoms and bonds.
Figure 15: Random valid samples from FMG trained on QM9. To the left of each molecule, the top 200 largest $\mathbf{u}_{a}$ and $\mathbf{u}_{b}$ values are shown along their corresponding extracted atoms and bonds.
F.2 GEOM-Drugs samples
Figures 16 and 17 depict a set of random valid samples from a model trained on GEOM.
Figure 16: Random valid samples from FMG trained on GEOM. To the left of each molecule, the top 200 largest $\mathbf{u}_{a}$ and $\mathbf{u}_{b}$ values are shown along their corresponding extracted atoms and bonds.
Figure 17: Random valid samples from FMG trained on GEOM. To the left of each molecule, the top 200 largest $\mathbf{u}_{a}$ and $\mathbf{u}_{b}$ values are shown along their corresponding extracted atoms and bonds.
Appendix G Classification-free guidance analysis
[…]
We determine how accurately the classification free guidance results in molecules containing the conditioned number of atoms. We compute the accuracy as the percentage of generated molecules in the correct conditioning bin $N$, the L1 distance between the binned number of atoms generated by the model $\tilde{N}$, and the conditioning bin $N$ as $\min(|\tilde{N}_{\text{high}}-N_{\text{low}}|,|N_{\text{high}}-\tilde{N}_{\text{low}}|)$ (or 0, when $N=\tilde{N}$), and “L1 lower” is the fraction of the total L1 distance attributed to incorrectly generated bins because of a strictly lower number of atoms generated $\tilde{N}<N$ (as a measure of distribution skewness).
@@ Appendix G Classification-free guidance analysis
Similar results in terms of accuracy and L1 distance can be noted in the GEOM-Drugs dataset (relative to the much larger possible number of atoms), but now the shift is in the opposite direction. This could be an artifact of the number of training iterations: early on in the QM9 dataset training experiments, we observed that the model tends to generate a lower number of atoms and bonds, which likely explains this opposite shift of the GEOM-Drugs method. For reference, our GEOM-Drugs model has been trained for about 6 epochs, compared to its QM9 counterpart (about 780).
Figure 18: Histogram of the generated number of atom classes $\tilde{N}$ and the conditioning used $N$. Small and more significant atom number distribution shifts occur for QM9 and GEOM-Drugs datasets, respectively. Together with the result of Table 3 which shows how critical it is to enforce the correct number of atoms for TVa, this result suggests that our method may remove atoms for stability, inducing an atom distribution shift.
@@ Appendix H Performance metric details
Percentage of valid molecules, computed using RDKit’s SanitizeMol function, computed using explicit H atoms.
@@
$\text{TV}_{\text{a}}=\sum_{a\in\mathcal{A}}\sum_{N_{a}=i}^{N}|p_{\mathcal{D}}(N_{a})-p_{g}(N_{a})|$
$\text{TV}_{\text{b}}=\sum_{b\in\mathcal{B}}\sum_{N_{b}=i}^{N}|p_{\mathcal{D}}(N_{b})-p_{g}(N_{b})|,$
where $N_{a},N_{b}$ are the numbers of atoms and bonds of type $b$, $a$ and $b$ are the atom and bond types, $\mathcal{A}$ and $\mathcal{B}$ are the sets of atoms and bonds considered, $p_{\mathcal{D}}$ and $p_{g}$ are probability distributions based on data and generated statistics, and ||$\cdot$|| ⋅ | is the absolute value.
@@
BL, BA, CEMMFF, and CExTB:
We determine the Wasserstein W1 distance between the generated and data bond angles, bond lengths, and Merck molecular force field energies (in kcal/mol).
The angle distributions are defined for all atom types that have two or more bonds, and we consider all pair-wise angles between position vectors $\mathbf{m}_{i}-\mathbf{m}_{k}$, and $\mathbf{m}_{j}-\mathbf{m}_{k}$, for any two atoms $i$ and $j$ connected to the same central atom $k$. We consider the average W1 distance over all atom types, and we denote it by BA.
@@
We use discretized bins of length $l_{\theta}=0.1$ degrees for angles, $l_{d}=0.01$Å for bond Euclidean distances, and $l_{e}=0.1$ (kcal/mol) for conformation energies, and approximate the generated and data cumulative distribution functions $F_{g}$ and $F_{d}$ and compute:
@@
$\text{BL}=l_{d}\frac{1}{N_{\mathcal{B}}}\sum_{b\in\mathcal{B}}\sum_{x\in B_{b}}|F_{d,b}(x)-F_{g,b}(x)|,$
where the bin upper limits $B_{a}$, $B_{b}$, and $B_{e}$ are chosen to contain all maximum and minimum values of the generated and training data samples, $N_{\mathcal{A}}$ and $N_{\mathcal{B}}$ are the number of unique atoms and bonds considered (e.g., we excluded H and F for BA computations, as these atoms never have two bonds).
For both Merck molecular force field (CEMMFF) and xTB (CExTB) we use the same bin length $l_{e}$ and compute the metics in the same way:
$\text{CE}=l_{e}\sum_{x\in B_{e}}|F_{d,e}(x)-F_{g,e}(x)|,$
where $F_{d,e}$ may be either MMFF or xTB CDF of generated and data energies.
Outlier problems of W1 distance:
Compared to our method, point-cloud methods may generate a molecule with an arbitrarily long bond length (i.e., $l\rightarrow\infty$), making the BL W1 metric arbitrarily high based on one or few bonds. Let $p_{\theta}$ and $p_{\mathcal{D}}$ be the generated and data bond length distributions. If their corresponding CDFs are $F_{\theta}$ and $F_{\mathcal{D}}$, the $W_{1}$ distance can be equivalently written as
[…]
where the integration is over the generated and data molecule lengths $l$. In fields, bond lengths are constrained by the chosen evaluated points $\mathcal{G}$, while in point clouds, $l$ may be arbitrarily high. In turn, this can theoretically incur arbitrarily high penalties due to a few outliers. In practice, this does not happen (Appendix A.1, Figure4, depicting the CDFs of bond lengths), and we further ensure we don’t have such an outlier advantage by restricting the bond lengths and angles to only those generated within the data limits when computing BA and EL metrics.